# The Fall of an Algorithm: Characterizing the Dynamics Toward Abandonment

### v1

@@ 4.6. Death
Overriding past decisions. Past decisions issued by an algorithm can continue to negatively impact people who received adverse outcomes. In the majority of cases of algorithms that were abandoned after their deployment, we found that the algorithm owners did not re-consider or override past (harmful) decisions made using the algorithm. For example, to our knowledge, rejected resumes were not re-screened (Dastin, 2018) and closed child welfare investigations were not re-opened (ap2, 2022). In a few notable cases, however, decisions were overturned and re-issued en masse. For example, in response to global protests against algorithm-assigned A-level grades, U.K. exam boards globally rescinded all algorithm-assigned grades and revised them based on teacher-assessed grades (without the algorithmic standardization under criticism (Ehsan et al., 2022)). Similarly, a legal settlement in Idaho guaranteed that people whose home care benefits were cut by an algorithm would continue to receive their original budgets (Hub, 2022b).
@@
Reincarnation. A growing number of investigations uncover how both successful and abandoned technologies can re-appear in different contexts (Kalluri et al., 2023; Loewenstein, 2023; Scott, 2023). We observed that when a city cancelled a contract with a technology vendor, the company often remained and continued to sell their product elsewhere (Haskins, 2019, 2021). Even as cities such as Palo Alto decided to drop their contracts with the predictive policing company PredPol, public records requests revealed that PredPol still had (many previously undisclosed) contracts with dozens of U.S. cities (Haskins, 2019). In response to bad press, some algorithm owners rebranded (sto, 2021): e.g., one month after the LAPD stopped using PredPol, the company changed its name to Geolitica and its slogan from “The Predictive Policing Company” to “Data-Driven Community Policing” (Bhuiyan, 2021). In 2023, SoundThinking (the company that created ShotSpotter, another prominent law enforcement algorithm) acquired IP from Geolitica, and still maintains versions of its predictive policing algorithms (Mehrotra and Cameron, 2023). The case of PredPol exemplifies how built technologies are valued as infrastructure that remains and can be recycled even after they are abandoned.
@@ 5. Factors relevant to algorithm abandonment
Users vs. Subjects. Is the algorithm a consumer product, or used on (unaware) subjects? Algorithms in consumer products (e.g., online platforms) ”fell” more quickly than algorithms used by organizations (e.g., governments) to analyze or make decisions about subjects who were potentially unaware of the algorithm’s existence. Direct users or consumers who interacted with technologies were more likely to be aware that there is an algorithm in place and have more agency to resist harmful interactions with it. In cases where the direct user who consumed the algorithm’s outputs used it to make decisions about someone else (e.g., a policing agency using facial recognition on city residents), impacted individuals were often unaware of its use or existence, which prolonged the process leading to abandonment. This distinction draws attention to the importance of understanding power differentials in who purchases and uses, versus who is most (negatively) impacted by algorithmic systems. Algorithm owners and critics alike should aim to empower and amplify the perspectives of those most impacted by algorithmic systems, even if they are not direct users.
@@
Value proposition. What is the ”value proposition” behind the use of the algorithm? A critical factor relevant to cases’ outcomes was understanding algorithm owners’ motivation for creating or adopting the algorithm, i.e., its value proposition. Publicly available value propositions often articulated a clear set of political values, e.g., to reduce MedicAid spending (by cutting care) (Weil and Edwards, 2023), to maximize returns to landlords (by raising rents) (Vogell et al., 2022), or to enable constant and wide-scale public surveillance (Dastin, 2020). In some cases, algorithm owners valued that automation allowed them to de-personalize decisions formerly made subjectively by humans, e.g., that leasing agents or nurse assessors had “too much empathy” (Vogell et al., 2022) to raise rents or cut hours. When an algorithm’s value proposition was clearly stated, in some cases critics organized successful campaigns that countered them (e.g., challenging the notion that surveillance makes us feel safe (log, 2020)). Algorithm owners should make an algorithm’s value proposition clear to the public. Critics should search for value propositions (i.e., how is the technology being marketed to its buyers? (Roemmich et al., 2023)) and make them visible so that they can be contested.
@@
The ”domino effect” caused by deep dependencies. How embedded is the algorithm into other systems? What other practices and processes ”depend” on it? Stand-alone algorithms released as independent products, like chatbots, are much easier to abandon than those embedded into larger social or technical systems, like algorithms used to automate important bureaucratic decisions (Alkhatib and Bernstein, 2019). Abandoning an algorithm that other components rely on might have a domino effect that requires changing its surrounding socio-technical components. For example, when courts ruled algorithms that assessed beneficiaries’ eligibility illegal, public officials scrambled to develop a new process (e.g., train human assessors) to allocate care resources (Hub, 2022a). Algorithm owners should aim to minimize dependencies by creating and budgeting for back-up plans for a scenario where the algorithm is no longer in use (a proposal that’s become popular in artificial intelligence governance (of Canada, 2023; of Science and Policy, 2022)).
@@
Procedural transparency and public notice. Is the public aware that an algorithm exists? Do we know how important decisions were made? We found that many harmful algorithms were used for months (or even years) before they were discovered and the general public was made aware of their existence. In contrast, when algorithm owners made their plans to use an algorithm public before it was deployed (like Missouri’s NF LOC algorithm (Hub, 2022c)), in some cases they decided to abandon the algorithm before it could inflict harm in the real world. Thus, the extent to which algorithm owners made information about the algorithm public, and when, played a pivotal role in the time it took for harmful systems to be abandoned. Algorithm owners should provide public notice and publicly release information about the algorithm’s design at the earliest possible stages of the algorithm’s lifecycle. FAccT researchers and advocacy groups can contribute to the small (yet growing) number of tools identified by Raji et al. (2024) designed to aid “target identification”, i.e., to “uncover deployed systems by making them visible to external stakeholders”. Finally, policymakers or organizations can institute transparency requirements for algorithm owners (e.g., those proposed in (Reisman et al., 2018; Moss et al., 2021; Gehami and Doty, 2023)).
@@
Access and auditability. What access do independent actors, e.g., members of the public, have to the algorithm? As discussed in Section 4.2, while independent audits were an extremely effective tool that triggered the decision to abandon an algorithm, independent actors had varying degrees of access to different algorithmic systems. FAccT researchers can contribute to infrastructure and build tools to support independent audits, such as those taxonomized by (Raji et al., 2024). One promising trend is that coalitions of different actors have started collaborations to share knowledge and data about their interactions with algorithmic systems, such as academic or non-profit research groups supporting public defenders and worker’s unions (Weil and Edwards, 2023; Hub, 2022c; FairFare, 2023). Policymakers and advocacy groups can consider funding these independent audits (Mozilla, 2023).
@@
Oversight and regulatory environment. Are policymakers receptive to creating new policies (within their political interests)? Who is held liable for algorithmic harms? As discussed in Section 4.5, instituting new “bright-line rules” (ain, 2023) that explicitly ban specific types of algorithms takes far less time (and material investment) compared to interpreting existing laws in court. Policymakers can consider instituting such bans proactively, even before the banned systems are in use, like the City of Somerville’s ban on employee’s use of facial recognition (Wu, 2019). Another trend we observed is a lack of accountability under the law for vendors that develop algorithmic systems and sell them to organizations that oversee their deployment. For example, when courts ruled Arkansas’ roll-out of an algorithm developed by interRAI (an independent vendor) illegal, the state of Arkansas, not interRAI, was required to pay the full settlement amount to plaintiffs who were harmed (Ly, 2023). This accountability structure means that even when one organization decides to stop using a harmful algorithm, it can continue to be used elsewhere with little consequence to the vendor who would continue to benefit from the algorithm (Lima et al., 2023). Policymakers should create and interpret existing laws to promote what Metcalf (2023) calls “distributed accountability”, where vendors are also held liable for remedying algorithmic harms.
@@ 6. Conclusion & Future Work
Grassroots movements, moratorium bans, and boardroom decisions to abandon harmful algorithmic systems continue to grow across the world. Yet, little academic attention has been paid to understanding the phenomena of algorithm abandonment. In our analysis of real-world cases involving (calls for) algorithm abandonment, we identified common phases of events that took place and key socio-technical factors that enabled (or prolonged) abandonment efforts. Our analysis reveals power imbalances and information asymmetries that prevent critics from holding algorithm owners accountable. Yet, victories from below in our case set illustrate how these power differentials might be challenged through policy interventions and collective action. In our discussion of relevant factors, we highlight steps that algorithm owners can take to create algorithms that can be more easily be contested and abandoned. Regulators can consider making these steps mandatory, such as requiring algorithm owners to make important information public at early stages of design and development (Reisman et al., 2018).
Our preliminary work points to several exciting directions for future work on algorithm abandonment. Researchers can take inspiration from successful collaborations that have built tooling and infrastructure to enable collective contestation and auditing, such as partnerships between academic researchers and worker’s unions (FairFare, 2023). Future research efforts can conduct more in-depth studies, such as interviews, to understand the experiences of actors involved in making the decision to abandon an algorithm, as well as other important stakeholders such as impacted communities, advocacy groups, and policymakers. Future work can also center campaigns that go beyond calls to abandon, and instead fight to abolish harmful technologies. In her conceptualization of “abolitionist technologies”, Ruha Benjamin writes that “calls for abolition are never simply about bringing harmful systems to an end but also about envisioning new ones. After all, the etymology of abolition includes Latin root words for ‘destroy’ (abolere) and ‘grow’ (olere)” (Benjamin, 2019). In our analysis, we found that algorithms in our case set implicate oppressive social systems or politics that informed their design. How can the FAccT community can better support the many advocacy initiatives (Benjamin, 2019; log, 2020; Milner, 2020; sto, 2021) that work to contest the values that animate harmful algorithmic systems, and imagine more just futures?
@@ 7. Impact Statements
Ethical Considerations. In proposing a process-oriented view of the dynamics of algorithm abandonment, our analysis reveals commonalities that place seemingly dissimilar cases (of chatbots (Hoover, 2023), B2B SAAS products (Vogell et al., 2022), and carceral technologies (System, 2021)) in conversation with one another. These similarities enable us to formulate recommendations for future research and policy interventions related to algorithm abandonment. However, our analysis has several limitations. In focusing on what is shared between a large and diverse set of algorithmic systems that are deployed in a wide variety of political contexts, we exclude important details that illustrate the uniqueness and complexity of each case. Further, given the rapidly evolving nature of algorithmic capabilities and harms, we acknowledge that our process-oriented view may fail to describe the events leading to the fall of all algorithmic systems (including recent generative, multi-modal, and multi-purpose AI systems). Because of these concerns, the 6Ds of abandonment are not meant to serve as a model of steps for critics to follow, but rather provide descriptive context to understand similar historic struggles.
@@ Acknowledgements.
- aia ([n. d.]) [n. d.]. AI, Algorithmic, and Automation Incidents and Controversies (AIAAIC) Repository. Retrieved January 18, 2024 from https://www.aiaaic.org/aiaaic-repository
- gal ([n. d.]) [n. d.]. Galactica Demo — galactica.org. https://galactica.org/. [Accessed 22-01-2024].
- was ([n. d.]) [n. d.]. New liver transplant rules yield winners, losers as wasted organs reach record high — washingtonpost.com. https://www.washingtonpost.com/business/2023/03/21/liver-transplants-acuity-circle-policy/. [Accessed 22-01-2024].
@@
- DeVrio et al. (2023) Alicia DeVrio, Motahhare Eslami, and Kenneth Holstein. 2023. Building, Shifting, & Employing Power: A Framework for Understanding & Empowering Action From Below in Response to Algorithmic Harm. In submission (2023).
@@
- Raji et al. (2024) Inioluwa Deborah Raji, Ryan Steed, Victor Ojewale, Briana Vecchione, , and Abeba Birhane. 2024. Towards Accountability Infrastructure: Gaps and opportunities in AI audit tooling. (2024). Preprint (Under Review).
@@
- Wong et al. (2021) Andrew Wong, Erkin Otles, John P. Donnelly, Andrew Krumm, Jeffrey McCullough, Olivia DeTroyer-Cooley, Justin Pestrue, Marie Phillips, Judy Konye, Carleen Penoza, Muhammad Ghous, and Karandeep Singh. 2021. External Validation of a Widely Implemented Proprietary Sepsis Prediction Model in Hospitalized Patients. JAMA Internal Medicine 181, 8 (08 2021), 1065–1070. https://doi.org/10.1001/jamainternmed.2021.2626 arXiv:https://jamanetwork.com/journals/jamainternalmedicine/articlepdf/2781307/jamainternal_wong_2021_oi_210027_1627674961.11707.pdf
- Wu (2019) Sarah Wu. 2019. Somerville City Council passes facial recognition ban. Retrieved January 16, 2024 from https://www.bostonglobe.com/metro/2019/06/27/somerville-city-council-passes-facial-recognition-ban/
@@
Appendix A Case List
| Case Name | Case Description | |
| Online platforms | ||
| 1 | Twitter Cropping Algorithm (Yee et al., 2021) | Users hypothesized that an algorithm that cropped images to show feed previews compatible with device aspect ratios noticed that it was more likely to crop out Black people (or other minority groups, i.e., women) in images with multiple people. |
| Chatbot | ||
| 4 | U.S. National Eating Disorder Assocation (NEDA)’s “Tessa” Wellness Bot (Wells, 2023) | Shortly NEDA fired all of their hotline staff (following their decision to unionize), eating disorder activists posted on social media about how Tessa (NEDA’s chatbot) gave harmful and triggering advice. |
| 5 | Facebook Galactica (gal, [n. d.]) | Within three days of releasing a public demo of their Galactica language model, Meta took down the public demo and re-branded Galactica as ”available for researchers” only. |
| 6 | South Korean Startup (Scatter Lab)’s ”Lee-Luda” Chatbot (https://www.theguardian.com/profile/justinmccurry, [n. d.]) | Seoul-based startup’s AI-chatbot ”Lee-Luda” was removed from Facebook messenger 20 days upon launch after users complained it was using offensive language, discriminatory and generated hate speech towards LGBTQ folks and people with disabilities. |
| Employment | ||
| 7 | Amazon AI Recuriting Tool (Dastin, [n. d.]) | Amazon disbanded an internal team whose goal was to train ML to screen resumes. |
| Child Welfare | ||
| 8 | Oregon DHS’s Safety at Screening Tool (ap2, 2022) | Oregon DHS decided to stop using an ”AI”-based screening tool (which is modeled after the AFST), in the wake of an April AP news article reporting that these tools have the ability to ”heighten racial disparities in the child welfare system”. |
| Case Name | Case Description | |
| Child Welfare | ||
| 11 | New Zealand’s ”Vulnerable Children PRM” (Predictive Risk Modeling) (Bukowitz and O’Loughlin, 2022) | New Zealand’s Social Development Minister intervenes and halts the deployment of the algorithm after seeing major concerns from additional research during the request for its ethics approval for further study. |
| Family & Social Services | ||
| 12 |
Teen Pregnancy Risk
Prediction in Argentina (Aplicada, 2018) |
| 13 |
The Spanish Ministry of the Interior’s “VioGén”
Gender Violence Recidivism Risk
Assessment (eti, 2022) |
The Eticas Foundation partnered with victims to publish an external audit after the Spanish Ministry declined requests of auditing the algorithm that underestimated level of risk in domestic and intimate partner violence cases leading to adverse outcomes. |
| Benefits | ||
| Case Name | Case Description | |
| Benefits | ||
| 18 |
Netherlands’ System Risk
Indication (SyRI) (van Veen, 2020) |
After local civil rights organizations filed a case against the state to stop algorithmic use for detecting fraud of welfare benefits targeting poor and marginalized neighborhoods, the Dutch court ruled that the algorithm violated the European Convention on Human Rights (ECHR) protections and halted it operation. |
| Healthcare | ||
| 19 | EPIC’ Sepsis Prediction Tool (Ross, 2021) | Epic, an American healthcare software company (and prominent EHR vendor), launched the ”Epic Sepsis Model” (ESM) in 2017, a proprietary LR model that alerts patients at high risk of sepsis from data ”routinely recorded within the EHR”. From 2021 onward, external and independent audits or investigations showed that the model was flawed; e.g., it had lower AUC than reported, had a high false alarm rate, etc. |
| 20 | US HHS and UNOS (the United Network for Organ Sharing)’s ”Acuity Circles Policy” (was, [n. d.]) | A joint investigation by the Washington Post and The Markup exposed how a new shift in the UNOS allocation algorithm’s criteria in removing distance as a factor for organ donation and prioritizing sickest patients regardless of location, drastically decreased the number of transplant surgeries and increased deaths in poorer states. |
| Case Name | Case Description | |
| Property Tech | ||
| 24 | RealPage’s YieldStar Algorithm (Vogell et al., 2022) | RealPage, a Texas-based company that provides property management software for landlords, developed an algorithm, ”YieldStar”, that suggests daily prices for open rental units. Critics claim that the software brings anti-trust concerns and has resulted in rent inflation across the U.S. |
| Education | ||
| 26 |
SAS Institute’s Educational Value-Added Assessment
System (EVAAS) in Houston (Capo and Bass, 2017) |
| Facial Recognition | ||
| 31 |
Facial Recognition in NYC
Atlantic Plaza Towers (Gagne, 2019) |
| Case Name | Case Description | |
| Facial Recognition | ||
| 32 | RiteAid’s Enrollement Database for Surveillance Systems (Dastin, 2020) | RiteAid, a large U.S. drugstore chain, added facial recognition surveillance systems in over 200 stores. Employees used the FR to surveil customers (e.g., ”persons of interest” who might attempt to engage in criminal activity). RiteAid stopped using FR after a Reuters investigation broke in 2020, and in 2023 were issued a five-year ban on their use of FR by the FTC. |
| 37 |
Uganda’s
”Intelligent Transport Monitoring System” (Nyeko, 2023) |
| Predictive Policing | ||
| 39 | City of Pittsburgh (in Pittsburgh, 2020a) | Pittsburgh City Council voted to ban ”obtaining, retaining, accessing, or using FR or predictive policing technology” after facing criticism from student activists for a Metro21 predictive policing experiment. |
| Case Name | Case Description | |
| Predictive Policing | ||
Appendix B Methodology (extended)

### v2

@@ 4.6. Death
Overriding past decisions. Past decisions issued by an algorithm can continue to negatively impact people who received adverse outcomes. In the majority of cases of algorithms that were abandoned after their deployment, we found that the algorithm owners did not re-consider or override past (harmful) decisions made using the algorithm. For example, to our knowledge, rejected resumes were not re-screened (Dastin, 2018) and closed child welfare investigations were not re-opened (ap2, 2022). In a few notable cases, however, decisions were overturned and re-issued en masse. For example, in response to global protests against algorithmically-assigned A-level grades, U.K. exam boards globally rescinded all algorithm-assigned grades and revised them based on teacher-assessed grades (without the algorithmic standardization under criticism (Ehsan et al., 2022)). Similarly, a legal settlement in Idaho guaranteed that people whose home care benefits were cut by an algorithm would continue to receive their original budgets (Hub, 2022b).
@@
Reincarnation. A growing number of investigations uncover how abandoned technologies can re-appear in different contexts (Kalluri et al., 2023; Loewenstein, 2023; Scott, 2023). We observed that when a city cancelled a contract with a technology vendor, the company often remained and continued to sell their product elsewhere (Haskins, 2019, 2021). Even as cities such as Palo Alto decided to drop their contracts with the predictive policing company PredPol, public records requests revealed that PredPol still had (many previously undisclosed) contracts with dozens of U.S. cities (Haskins, 2019). In response to bad press, some algorithm owners rebranded (sto, 2021): e.g., one month after the LAPD stopped using PredPol, the company changed its name to Geolitica and its slogan from “The Predictive Policing Company” to “Data-Driven Community Policing” (Bhuiyan, 2021). In 2023, SoundThinking (the company that created ShotSpotter, another prominent law enforcement algorithm) acquired IP from Geolitica, and still maintains versions of its predictive policing algorithms (Mehrotra and Cameron, 2023). The case of PredPol exemplifies how built technologies are valued as infrastructure that remains and can be recycled even after they are abandoned.
@@ 5. Factors relevant to algorithm abandonment
Users vs. Subjects. Is the algorithm a consumer product, or used on (unaware) subjects? Algorithms in consumer products (e.g., online platforms) ”fell” more quickly than algorithms used by organizations (e.g., governments) to analyze or make decisions about subjects who were potentially unaware of the algorithm’s existence. Direct users or consumers who interacted with technologies were more likely to be aware that there was an algorithm in place, and had more agency to resist harmful interactions with it. In cases where the direct user who consumed the algorithm’s outputs used it to make decisions about someone else (e.g., a policing agency using facial recognition on city residents), impacted individuals were often unaware of its use or existence, which prolonged the process leading to abandonment. This distinction draws attention to the importance of understanding power differentials in who purchases and uses, versus who is most (negatively) impacted by algorithmic systems. Algorithm owners and critics alike should aim to empower and amplify the perspectives of those most impacted by algorithmic systems, even if they are not direct users.
@@
Value proposition. What is the ”value proposition” behind the use of the algorithm? A critical factor relevant to cases’ outcomes was understanding algorithm owners’ motivation for creating or adopting the algorithm, i.e., its value proposition. Publicly available value propositions often articulated a clear set of political values, e.g., to reduce Medicaid spending (by cutting care) (Weil and Edwards, 2023), to maximize returns to landlords (by raising rents) (Vogell et al., 2022), or to enable constant and wide-scale surveillance (Dastin, 2020). In some cases, algorithm owners valued that automation allowed them to de-personalize decisions formerly made by humans, e.g., that leasing agents or nurse assessors had “too much empathy” (Vogell et al., 2022) to raise rents or cut hours. When an algorithm’s value proposition was clearly stated, in some cases critics organized successful campaigns that countered them (e.g., challenging the notion that surveillance makes us feel safe (log, 2020)). Algorithm owners should make an algorithm’s value proposition clear to the public. Critics should search for value propositions (i.e., how is the technology being marketed to its buyers? (Roemmich et al., 2023)) and make them visible so that they can be contested.
@@
The ”domino effect” caused by deep dependencies. How embedded is the algorithm into other systems? What other practices and processes ”depend” on it? Stand-alone algorithms released as independent products, like chatbots, were much easier to abandon than those embedded into larger social or technical systems, like algorithms used to automate important bureaucratic decisions (Alkhatib and Bernstein, 2019). Abandoning an algorithm that other components rely on might have a domino effect that requires changing its surrounding socio-technical components. For example, when courts ruled algorithms that assessed beneficiaries’ eligibility illegal, public officials scrambled to develop a new process (e.g., to train human assessors) to allocate care resources (Hub, 2022a). Algorithm owners should aim to minimize dependencies by creating and budgeting for back-up plans in a scenario where the algorithm is no longer in use, a proposal that has become popular in recent AI governance frameworks (of Canada, 2023; of Science and Policy, 2022).
@@
Procedural transparency and public notice. Is the public aware that an algorithm exists? Do we know how important decisions were made? We found that many harmful algorithms were used for months (or even years) before they were discovered and the general public was made aware of their existence. In contrast, when algorithm owners made their plans to use an algorithm public before it was deployed (like Missouri’s NF LOC algorithm (Hub, 2022c)), in some cases they decided to abandon the algorithm before it could inflict harm in the real world. Thus, the extent to which algorithm owners made information about the algorithm public, and when, played a pivotal role in the time it took for harmful systems to be abandoned. Algorithm owners should provide public notice and publicly release information about the algorithm’s design at the earliest possible stages of the algorithm’s lifecycle. FAccT researchers and advocacy groups can contribute to the small (yet growing) number of tools identified by Ojewale et al. (2024) designed to aid “target identification”, i.e., to “uncover deployed systems by making them visible to external stakeholders”. Finally, policymakers or organizations can institute transparency requirements for algorithm owners (e.g., those proposed in (Reisman et al., 2018; Moss et al., 2021; Gehami and Doty, 2023)).
@@
Access and auditability. What access do independent actors, e.g., members of the public, have to the algorithm? As discussed in Section 4.2, while independent audits were an extremely effective tool that triggered the decision to abandon an algorithm, independent actors had varying degrees of access to different algorithmic systems. FAccT researchers can contribute to infrastructure and build tools to support independent audits, such as those taxonomized by (Ojewale et al., 2024). One promising trend is that coalitions of different actors have started collaborations to share knowledge and data about their interactions with algorithmic systems, such as academic or non-profit research groups supporting public defenders and worker’s unions (Weil and Edwards, 2023; Hub, 2022c; FairFare, 2023). Policymakers and advocacy groups can consider funding these independent audits (Mozilla, 2023).
@@
Oversight and regulatory environment. Are policymakers receptive to creating new policies? Who is held liable for algorithmic harms? As discussed in Section 4.5, instituting new “bright-line rules” (ain, 2023) that explicitly ban specific types of algorithms takes far less time (and material investment) compared to interpreting existing laws in court. Policymakers can consider instituting such bans proactively, even before the banned systems are in use, like the City of Somerville’s ban on employees’ use of facial recognition (Wu, 2019). Another trend we observed is a lack of accountability under the law for vendors that develop algorithmic systems and sell them to organizations that oversee their deployment. For example, when courts ruled Arkansas’ roll-out of an algorithm developed by interRAI (an independent vendor) illegal, the state of Arkansas, not interRAI, was required to pay the full settlement amount to plaintiffs who were harmed (Ly, 2023). This accountability structure means that even when one organization decides to stop using a harmful algorithm, it can continue to be used elsewhere with little consequence to the vendor who would continue to benefit from the algorithm (Lima et al., 2023). Policymakers should create and interpret existing laws to promote what Metcalf (2023) calls “distributed accountability”, where vendors are also held liable for remedying algorithmic harms.
@@ 6. Conclusion & Future Work
Grassroots movements, moratorium bans, and boardroom decisions to abandon harmful algorithmic systems continue to grow across the world. Yet, comparatively little academic attention has been paid to understanding the phenomena of algorithm abandonment. In our analysis of real-world cases involving (calls for) algorithm abandonment, we identified common phases of events that took place and key socio-technical factors that enabled (or prolonged) abandonment efforts. Our analysis reveals power imbalances and information asymmetries that prevent critics from holding algorithm owners accountable. Yet, victories from below in our case set illustrate how these power differentials might be challenged through policy interventions and collective action. In our discussion of relevant factors, we highlight steps that algorithm owners can take to create algorithms that can be more easily be contested and abandoned. Regulators can consider making these steps mandatory, such as requiring algorithm owners to make important information public at early stages of design and development (Reisman et al., 2018).
Our preliminary work points to several exciting directions for future work on algorithm abandonment. Researchers can take inspiration from successful collaborations that have built tooling and infrastructure to enable collective contestation, such as partnerships between academic researchers and worker’s unions (Weil and Edwards, 2023; Hub, 2022c; FairFare, 2023). Future research efforts can conduct more in-depth studies, such as interviews, to understand the experiences of actors involved in making the decision to abandon an algorithm, as well as other important stakeholders such as impacted communities, advocacy groups, and policymakers. Future work can also center campaigns that go beyond calls to abandon, and instead fight to abolish harmful technologies. In her conceptualization of “abolitionist technologies”, Ruha Benjamin writes that “calls for abolition are never simply about bringing harmful systems to an end but also about envisioning new ones. After all, the etymology of abolition includes Latin root words for ‘destroy’ (abolere) and ‘grow’ (olere)” (Benjamin, 2019). In our analysis, we found that algorithms in our case set implicate oppressive social systems or politics that informed their design. How can the FAccT community can better support the many advocacy initiatives (Benjamin, 2019; log, 2020; Milner, 2020; sto, 2021) that work to contest the values that animate harmful algorithmic systems, and imagine more just futures?
@@ 7. Impact Statements
Ethical Considerations. In proposing a process-oriented view of the dynamics of algorithm abandonment, our analysis reveals commonalities that place seemingly dissimilar cases (of chatbots (Hoover, 2023), B2B SAAS products (Vogell et al., 2022), and carceral technologies (System, 2021)) in conversation with one another. These similarities enable us to formulate recommendations for future research and policy interventions related to algorithm abandonment.
However, our analysis has several limitations. In focusing on what is shared between a large and diverse set of algorithmic systems that are deployed in a wide variety of political contexts, we exclude important details that illustrate the uniqueness and complexity of each case. Further, given the rapidly evolving nature of algorithmic capabilities and harms, we acknowledge that our process-oriented view may fail to describe the events leading to the fall of all algorithmic systems (including recent generative, multi-modal, and multi-purpose AI systems). Because of these concerns, the 6 D’s of abandonment are not meant to serve as a model of steps for critics to follow, but rather provide descriptive context to understand similar historic struggles.
@@ Acknowledgements.
- gal (2022) 2022. Galactica Demo — galactica.org. https://galactica.org/. [Accessed 22-01-2024].
@@
- was (2023) 2023. New liver transplant rules yield winners, losers as wasted organs reach record high — washingtonpost.com. https://www.washingtonpost.com/business/2023/03/21/liver-transplants-acuity-circle-policy/. [Accessed 22-01-2024].
@@
- AIAAIC (2023) AIAAIC. 2023. AI, Algorithmic, and Automation Incidents and Controversies (AIAAIC) Repository. Retrieved January 18, 2024 from https://www.aiaaic.org/aiaaic-repository
@@
- DeVrio et al. (2023) Alicia DeVrio, Motahhare Eslami, and Kenneth Holstein. 2023. Building, Shifting, & Employing Power: A Taxonomy of Responses From Below to Algorithmic Harm. In submission (2023).
@@
- Ojewale et al. (2024) Victor Ojewale, Ryan Steed, Briana Vecchione, Abeba Birhane, and Inioluwa Deborah Raji. 2024. Towards AI Accountability Infrastructure: Gaps and Opportunities in AI Audit Tooling. arXiv:2402.17861 [cs.CY]
@@
Wong et al. (2021)
Andrew Wong, Erkin Otles,
John P. Donnelly, Andrew Krumm,
Jeffrey McCullough, Olivia
DeTroyer-Cooley, Justin Pestrue, Marie
Phillips, Judy Konye, Carleen Penoza,
Muhammad Ghous, and Karandeep Singh.
2021.
External Validation of a Widely Implemented
Proprietary Sepsis Prediction Model in Hospitalized Patients.
JAMA Internal Medicine
181, 8 (08
2021), 1065–1070.
https://doi.org/10.1001/jamainternmed.2021.2626
arXiv:https://jamanetwork.com/journals/jamainternalmedicine/articlepdf
/2781307/ jamainternal_wong_2021_oi_210027_1627674961.11707.pdf - Wu (2019) Sarah Wu. 2019. Somerville City Council passes facial recognition ban. Retrieved January 16, 2024 from https://www.bostonglobe.com/metro/2019/06/27/somerville-city-council-passes-facial-recognition-ban/
@@
Appendix A Methodology (extended)